OpenStudy (anonymous):

The perimeter of a kite is 52 meters. If one side measure 19 meters find the measure of the other side.

OpenStudy (anonymous):

The adjacent side would also be 19. So to find the other side you would want to add 19 + 19 which equals 38. Subtract it from 52. You now have 14. But since their are two more sides divide that by two. 

OpenStudy (anonymous):

7 :D Thanks :P

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

Can you do anymore?

OpenStudy (anonymous):

I could help with a couple more.

OpenStudy (anonymous):

This one is troubling...

OpenStudy (anonymous):

hmm.. this one I dont remember how to solve.

OpenStudy (anonymous):

I can help. We can apply the same method of the kite problem you were looking at earlier. In this problem, we are given a kite as well, but the lengths are expressed as expressions instead of numbers. So, we are given the perimeter of the kite as 36 ft. An easy way to solve this problem is to lump all of the expressions into the same part and solve for x.(A terrible way to explain I know but bear with me) So what we must do is take all the factors and add them together to equal the perimeter. We get: 2(8x-3)+2(2x+1)=36 16x-6+4x+2=36 20x-4=36 20x=40 x=2 We get the 2 in front of the 8x-3 and also in front of the 2x+1 because we have 2 sides of each of those expressions if that makes sense... Hope it makes sense to you now!